OpenStudy (anonymous):

\[\sum_{K=1}^{\infty} 4(\frac{ 1 }{ 2 })^{k-1}\]

OpenStudy (anonymous):

Evaluate the following

OpenStudy (jhannybean):

This is a geometric series, right?

OpenStudy (anonymous):

Take 4 outside. It becomes an infinite GP with common ratio 1/2

OpenStudy (zzr0ck3r):

what is the common ratio?

OpenStudy (zzr0ck3r):

what is the first term?, what do we do when we know these two things?

OpenStudy (anonymous):

First term is 1

OpenStudy (anonymous):

wait i thought first term was 4?

OpenStudy (jhannybean):

yes, it is.

OpenStudy (anonymous):

Common Ratio 1/2... Calculate sum And multiply the sum by 4 ! 4 is isolated.

OpenStudy (luigi0210):

4(1/2)^0 4(1) =4

OpenStudy (zzr0ck3r):

factor out the 4

OpenStudy (jhannybean):

\[\large \sum_{n=1}^{\infty}a_{1}r^{n-1}\] so a1 =4.

OpenStudy (anonymous):

answer would be 8 !! (Final Answer)

OpenStudy (anonymous):

The answer is supposed to be 8 but i dont quite get how you get taht

OpenStudy (anonymous):

Factor out 4.. and forget it Calculate Sum 1 + 1/2 + 1/4 + 1/8 ...... upto infinity This sum is 2.. Multiply by 4 and you get 8

OpenStudy (jhannybean):

wait...by "evaluating" are you looking to see if it's convergent/divergent, and if it's convergent,finding the sum?

OpenStudy (zzr0ck3r):

factor out the 4, then treat it as (1/2)^n and you have first term is 1 then sum is 1/(1-1/2) = 2 then multiply by 4 and get 8

OpenStudy (anonymous):

It's obviously convergent..

OpenStudy (anonymous):

pretty sure its convergent

OpenStudy (zzr0ck3r):

it is, because (1/2) <1

OpenStudy (anonymous):

so i use the sum of geometric formula then i multiply by 4?

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

YEs ! @Yellowpanda

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

welcome :)

OpenStudy (zzr0ck3r):

treat it like \[\sum{}{}(\frac{1}{2})^{n}\]

OpenStudy (zzr0ck3r):

then multiply by 4 because of ab+ac+ad....=a(b+c+d....

OpenStudy (jhannybean):

Ahh, i see. a1=4, r = 1/2 \[\large S_{n}= \frac{a}{1-r} = \frac{4}{1-(1/2)} = 8\]

OpenStudy (jhannybean):

And this only applies because the series is convergent....

OpenStudy (zzr0ck3r):

yes, and note, you could have just factored the 4 out, treated the first term as 1, then multiply the 4 back when you are done. That is what we were saying.

OpenStudy (zzr0ck3r):

some people get confused on why the ratio is (1/2) when there is a 4 next to it, so I factor it out to show it does not matter.

OpenStudy (jhannybean):

Ohh I see, I see.